OpenStudy (anonymous):

Given the parent functions f(x) = 5x − 1 and g(x) = 3^x − 9, what is g(x) − f(x)?

OpenStudy (anonymous):

wait im solving it

OpenStudy (anonymous):

g(x) − f(x) = 3^x − 8 − 5x g(x) − f(x) = 3^x − 10 − 5x g(x) − f(x) = −2x − 8 − 3x g(x) − f(x) = −2x − 10 − 3x

OpenStudy (anonymous):

option 1 is correct

OpenStudy (anonymous):

g(x) − f(x) =(3^x − 9)-(5x − 1 ) =3^x-9-5x+1 =3^x-5x-9+1 =3^x-5x-8 Option 1 is correct

OpenStudy (anonymous):

@julithehero I'm sure its option 1

OpenStudy (anonymous):

If f(x) = log3 (x + 1), what is f^−1(2)?

OpenStudy (anonymous):

its option a dont worry